OpenStudy (anonymous):

Find the x-intercept of -2x +y=4

OpenStudy (anonymous):

put y = 0 in this equation to get x-intercept

OpenStudy (anonymous):

That didn't help

OpenStudy (anonymous):

y=mx+b m=2, b=4 where m is the slope and b is the y-intercept

OpenStudy (anonymous):

you have this equation -2x +y=4 now put y = 0 in this equation you will get . . . do it

OpenStudy (anonymous):

Some of my answers are (-2,0),(0,-2),(0,4), (4,0)

OpenStudy (anonymous):

y not you try it instead of having direct answers

OpenStudy (anonymous):

I did but I don't understand it and I need help

OpenStudy (the_fizicx99):

To find the x intercept set y to 0. To find the y intercept set x for 0

OpenStudy (the_fizicx99):

-2x + y = 4 So you want to find x, set y to 0

OpenStudy (the_fizicx99):

What would that look like @Punkchick

OpenStudy (anonymous):

Would be (0,4)

OpenStudy (anonymous):

or would it be (0,-2)

OpenStudy (the_fizicx99):

No. Don't skip steps, -2x + y = 4 Set y for 0 and rewrite it.

OpenStudy (the_fizicx99):

An order pair is in the form of (x,y) .....

OpenStudy (anonymous):

So its (-2,0)

OpenStudy (the_fizicx99):

Yes. But how did you get that?

OpenStudy (anonymous):

I guess to be honest and can u help me o a couple more

OpenStudy (the_fizicx99):

You can do this lol, set y for 0. -2x + 0 = 4 Now to ISOLATE x you divide both sides by -2x So, 4/-2 = -2 x = -2 (-2,0)